A simple pendulum of length 0.4m has a period of 2.4s. The period of a simple pendulum of length 0.36 is
1
Expert's answer
2019-01-14T09:07:42-0500

The time period of a simple pendulum for small amplitude of oscillations depends on the length

of the pendulum

"L"

and the acceleration due to gravity

"g"

, and can be found as follows:

"T = 2 \pi \sqrt{\frac{L}{g}} = 2 \pi \sqrt{\frac{0.36 m}{9.8 \frac{m}{s^2}}} = 1.2 s."

Answer:

"T = 1.2 s"

.